# Spatial super-resolution in nanosensing with blinking emitters

### v1

@@ II.4 Reconstruction of parameter distribution for noisy signal
While for ideal noiseless data the constructed estimator (26) would provide reasonable local values for the investigated parameter distribution $\theta(\mathbf{r})$, its application to the noisy data (especially for sparsely located emitters) deserves additional attention. The structure of expression (22) implies that the cumulant values $C_{j}^{(n)}(\mathbf{r})$ estimated from finite-sample data are more reliable near emitters ($\mathbf{r}\approx\mathbf{r}_{k}$ for certain $k$), but may contain significant relative shot noise contribution in the region between the emitters where the signals are weak. The effect becomes more pronounced for higher-order cumulants, since the effective PSF $U^{n}(\mathbf{r}-\mathbf{r}_{k})$ is narrower for larger $n$, while the contribution of shot noise is generally growing for larger cumulant orders.
The inaccuracy $\Delta\bar{\theta}(\mathbf{r})$ of the reconstructed parameter value $\bar{\theta}(\mathbf{r})=\bar{\theta}_{\text{nonlin}}^{(n)}(\mathbf{r})$ is determined by the relative contribution of the shot noise to the quantities $C_{0}^{(n)}(\mathbf{r})$ and $C_{1}^{(n)}(\mathbf{r})$, which is, roughly, inversely proportional to the square root of the detected signal:
$\Delta\bar{\theta}(\mathbf{r})\propto\frac{1}{\sqrt{\left|C_{0}^{(n)}(\mathbf{r})+C_{1}^{(n)}(\mathbf{r})\right|}}.$ (27)
The likelihood of getting the estimated distribution $\bar{\theta}(\mathbf{r})$ for the true distribution $\theta(\mathbf{r})$ can be represented as
$L\left[\bar{\theta}(\mathbf{r})|\theta(\mathbf{r})\right]\propto\exp\left(-\int d^{2}\mathbf{r}\frac{[\bar{\theta}(\mathbf{r})-\theta(\mathbf{r})]^{2}}{2\Delta\bar{\theta}^{2}(\mathbf{r})}\right).$ (28)
@@
According to Bayes’ formula [41, 42], the posterior probability of the true distribution $\theta(\mathbf{r})$ conditioned by the estimated distribution $\bar{\theta}(\mathbf{r})$ equals
$p\left[\theta(\mathbf{r})|\bar{\theta}(\mathbf{r})\right]\propto L\left[\bar{\theta}(\mathbf{r})|\theta(\mathbf{r})\right]p_{0}[\theta(\mathbf{r})].$ (30)
@@
$\theta(\mathbf{r})=\operatorname*{argmax}\limits_{\theta(\mathbf{r})}p\left[\theta(\mathbf{r})|\bar{\theta}(\mathbf{r})\right].$ (31)
Practically, the parameter values are estimated at a discrete grid of points $\mathbf{r}_{ij}$. In that case, the derivatives in Eq. (30) can be replaced by finite differences, and the values $\theta(\mathbf{r}_{ij})$ can be found by unconstrained quadratic minimization (Appendix D).
@@ II.5 Sensing method summary
2. Sensing calibration. The optimal weights $Q_{0}(x)$ and $Q_{1}(x)$ of the two detection channels are defined in Appendix A. If the experimental capabilities do not allow to perform such optimal measurement, the actual weights $Q_{0}(x)$ and $Q_{1}(x)$ are defined by the available detection setup (the profiles of the channels’ sensitivity). For sensing based on the $n$-th order cumulant, the linear estimator is defined according to Eq. (24). The nonlinear relation $\bar{\theta}^{(n)}=\Theta_{n}(\theta)$ between the estimator and the true value of the parameter is found by theoretical consideration or direct measurements of the sample response for a uniform distribution of the parameter $\theta(\mathbf{r})=const$. The final nonlinear estimator is constructed according to Eq. (26).
@@
4. Data processing. For each channel $k$ and each detection position $\mathbf{r}_{ij}$ separately, the cumulant value $C_{k}^{(n)}(\mathbf{r}_{ij})=C_{t}^{(n)}[F_{k}(\mathbf{r}_{ij},t)]$ is calculated. For $n=2$, 3, 4 the expressions are provided by Eqs. (5)–(7) with the expectation values replaced by finite-sample averages. The local estimates of the parameter $\bar{\theta}_{\text{nonlin}}^{(n)}(\mathbf{r})$ are calculated according to Eq. (26).
@@
5. Data interpolation. To account for the difference of the estimate reliability at different points, the corrected parameter distribution is constructed according to Eq. (31) with the method described in Appendix D.
III Model experiment
III.1 Setup description
@@
To verify the proposed technique, we performed the following model experiment. We used a paper screen as a model of fluorescent sample. The screen was irradiated with 13 KY-008 lasers controlled by the microcontroller board Arduino MEGA [43]. The image of the screen was recorded with Thorlabs DCC1545M-GL camera with 50 mm lens (Fig. 2). To simulate the diffraction limit, the lens was placed 2 mm closer to the camera than it was necessary to get a clear picture. The images of clear and unfocused laser spots are shown in Fig. 3. Each laser played the role of a fluorescent particle. Correspondingly, each particle had prescribed “relaxation time” meaning the time between “excitation pulse” and “fluorescent emission”. To denote “excitation pulse” in the recording data, Arduino made all lasers blink two times. Since the minimal distinguishable time gap is 1 frame (about 43 ms), we used frames as a measure of time. Therefore, the “excitation pulse” was a series of 2 one-frame flashes of lasers separated by 1 frame of darkness. The “fluorescent emission” was 1-frame flashing of a laser. To simulate the fluorescent particles with different temperature, the lasers were divided into “hot” and “cold” ones. For the hot lasers, the “relaxation time” was 5 frames with standard deviation of 2 frames, while the cold lasers had “relaxation time” of 10 frames with the same standard deviation.
@@
Figure 2: The scheme of model experiment. The lasers controlled with Arduino microcontroller board irradiate the paper screen. The light scattered by the screen is collected with lens to the Thorlabs camera.
@@
Figure 3: The images of laser spots on the paper screen: (a) the lens is placed to get a clear image; (b) the lens is 2 mm closer to the camera than it is necessary for a clear image.
III.2 Data acquisition
@@
The data were recorded in series of 10000 frames with the frame rate of 24 fps. For each “excitation cycle”, each laser was randomly assigned its state (active or dark) and the time delay between the “excitation pulse” and its relaxation (“fluorescent emission”). The delay was sampled according to normal distribution with the mean determined by the model temperature of the laser (“hot” or “cold”). The random sampling was performed by the Arduino microcontroller. Then, Arduino produced an “excitation pulse” and made each active laser to flash after the time delay assigned to it. After modeling “fluorescent emission” of all the active lasers withing the current cycle (when all the lasers are switched off), Arduino waited for 1 frame, and repeated the cycle. In the end, we got a video containing intensity distribution in time and space.
III.3 Data processing and results
@@
First, the recorded video was divided into frames with ffmpeg program [44] and processed with Python script. The script calculated total intensity for each frame and identified “excitation pulses” according to specific alternation of high and low intensity of frames. The frames designating the “excitation pulse” were excluded from further calculations. According to the notations, introduced in Section II, the “excitation cycles” are indexed by $t=1$, …, $N_{\text{f}}$.
@@
The signal for each “excitation cycle” was split into two channels and accumulated pixelwise. The “hot” signal $I_{0}(i,j;t)$ is the pixelwise sum of intensities for the 1st - 7th frames after the $t$-th “excitation pulse”, where indices $i$ and $j$ enumerate pixel positions in each frame. The “cold” signal $I_{1}(i,j;t)$ is defined in the same way starting from the 8th frame after the “excitation pulse” (Fig. 4).
@@
Figure 4: Example of ”excitation cycle” timeline, where two pulses appear in the “hot” time window and two pulses - in the “cold” time window.
@@
The intensity (mean signal value) and the 2nd, 3rd, and 4th order cumulants for the two channels where defined according to Eqs. (5)–(8). The obtained intensity and cumulant images are shown in Fig. 5. For visualization purposes, the “cold” and “hot” channels were assigned to the blue and red channels of RGB palette respectively. Since the cumulant of the 3rd order can be negative, the absolute value of the cumulant was used to create the image. One can see that the intensity distribution (Fig. 5(a)) is blurred like original unfocused image (Fig. 3(b)). However, it is possible to distinguish the light of “hot” and “cold” lasers because of different relaxation times. The light spots are more visible in the second-order cumulant image (Fig. 5(b)), but the background caused by blurring is still strong. The third-order cumulant image (Fig. 5(c)) is clear enough and the background is hardly visible. The best quality image is given by the fourth-order cumulant (Fig. 5(d)). It allows one to distinguish clearly spots of “hot” and “cold” lasers. To summarize, the experiment is in a good agreement with the theory, and the calculation of cumulants of intensity fluctuations allows one to visualize the particles with different relaxation times.
@@
III.4 Reconstruction of local “temperature”
@@
To fit the above-described model experiment into the theoretical framework, discussed in Section II, we assign the normalized dimensionless “temperature” equal to $\theta=-1$ to “cold” lasers and $\theta=+1$ to “hot” lasers. The emission profile variable $x$ corresponds to the camera frame number within the “hyperframe” (excitation cycle) of SOFI in the considered case. For such definitions, the model emission profile is described as
$\varepsilon(x,\theta)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{(x-7.5+2.5\theta)^{2}}{2\sigma^{2}}\right)+b(x),$ (32)
where $\sigma=2$ and the additional parameter-independent background signal is modeled as $b(x)=\eta\exp\{-(x-7.5)^{2}/(2\sigma_{\text{b}}^{2})\}/(\sqrt{2\pi}\sigma_{\text{b}})$ with $\sigma_{\text{b}}=5$ and $\eta=0.1$.
A simple threshold filter is used to split the two detection channels:
$Y_{0}=\sum_{x\leq 7}\varepsilon(x,\theta),$ (33)
$Y_{1}=\sum_{x\geq 8}\varepsilon(x,\theta).$ (34)
The decomposition coefficients $A_{ij}$ (Eq. (17)) can be approximated as
$A_{j0}=\frac{1+\eta}{2},\quad A_{j1}=\pm\frac{1}{\sqrt{2\pi}},$ (35)
where $j=0$ and 1 correspond to the “$+$” and “--” signs respectively. The linear estimator $\bar{\theta}^{(n)}(\mathbf{r})$ is calculated according to Eq. (24) with $\theta_{0}=0$, while its more accurate nonlinear counterpart is constructed according to Eq. (26) using the procedure outlined before it (Fig. 6). Further, the reconstructed distribution can be refined using the method outlined in the Section II.4 and Appendix D.
Figure 6: The nonlinear estimator $\bar{\theta}_{\text{nonlin}}^{(n)}=\Theta^{-1}(\bar{\theta}^{(n)})$ for the model experiment and analysis of intensity ($n=1$) and cumulants ($n=2$, 3, 4). The dashed diagonal line is shown for visual guidance.
@@
The results are presented in Fig. 7. As predicted, usage of higher-order cumulants enhances spatial resolutions and enables analysis of finer features of the sample. According to the standard performance of SOFI, using fourth order cumulant instead of intensity leads to approximately twice narrower efficient PSF.
Figure 7: Reconstructed distribution of the model normalized temperature $\theta$ for analysis of intensity (a), the second (b), the third (c) and the fourth (d) order cumulants. Only the central part of the image is shown.
IV Discussion
To get quantitative estimates of the proposed approach performance, we applied several sharpness measures to the reconstructed distribution of the local “temperature”. The results shown in Fig. 8 indicate that the most blurred distribution is obtained for the traditional intensity-based approach. Derivative-based measures (Tenengrad and high-pass filtering [45]) give preference to the second-order cumulant (after application of noise-reduction algorithm from subsection II.4), while eigenvalue-based [46] and histogram entropy based [45] methods assign the best score to the fourth-order cumulant.
Figure 8: Sharpness of the reconstructed distribution of the model normalized temperature $\theta$ for different orders of cumulants used for the analysis. Solid (dashed) lines correspond to the distributions after (before) the noise correction described in subsection II.4.

### v2

@@ II.4 Reconstruction of parameter distribution for noisy signal
While for ideal noiseless data the constructed estimator (26) would provide reasonable local values for the investigated parameter distribution $\theta(\mathbf{r})$, its application to noisy data (especially for sparsely located emitters) deserves additional attention. The structure of expression (24) implies that the cumulant values $C_{j}^{(n)}(\mathbf{r})$ estimated from finite-sample data are more reliable near emitters ($\mathbf{r}\approx\mathbf{r}_{k}$ for certain $k$), but may contain significant relative shot noise contribution in the region between the emitters where the signals are weak. The effect becomes more pronounced for higher-order cumulants, since the effective PSF $U^{n}(\mathbf{r}-\mathbf{r}_{k})$ is narrower for larger $n$, while the contribution of shot noise is generally growing for larger cumulant orders.
The inaccuracy $\Delta\bar{\theta}^{(n)}(\mathbf{r})$ of the reconstructed parameter value $\bar{\theta}^{(n)}(\mathbf{r})$ is determined by the relative contribution of the shot noise to the quantities $C_{0}^{(n)}(\mathbf{r})$ and $C_{1}^{(n)}(\mathbf{r})$, which is, roughly, inversely proportional to the square root of the detected signal:
$\Delta\bar{\theta}^{(n)}(\mathbf{r})\propto\frac{1}{\sqrt{\left|C_{0}^{(n)}(\mathbf{r})+C_{1}^{(n)}(\mathbf{r})\right|}}.$ (30)
The likelihood of getting the estimated distribution $\bar{\theta}^{(n)}(\mathbf{r})$ for the true distribution $\theta(\mathbf{r})$ can be represented as
$L\bigl{[}\bar{\theta}^{(n)}(\mathbf{r})\big{|}\theta(\mathbf{r})\bigr{]}\propto\exp\left(-\int d^{2}\mathbf{r}\frac{\left[\bar{\theta}^{(n)}(\mathbf{r})-\theta(\mathbf{r})\right]^{2}}{2\left[\Delta\bar{\theta}^{(n)}(\mathbf{r})\right]^{2}}\right).$ (31)
@@
According to Bayes’ formula [70, 71], the posterior probability of the true distribution $\theta(\mathbf{r})$ conditioned by the estimated distribution $\bar{\theta}^{(n)}(\mathbf{r})$ equals
$p\bigl{[}\theta(\mathbf{r})\big{|}\bar{\theta}^{(n)}(\mathbf{r})\bigr{]}\propto L\bigl{[}\bar{\theta}^{(n)}(\mathbf{r})\big{|}\theta(\mathbf{r})\bigr{]}p_{0}[\theta(\mathbf{r})].$ (33)
@@
$\theta(\mathbf{r})=\operatorname*{argmax}\limits_{\theta(\mathbf{r})}p\bigl{[}\theta(\mathbf{r})\big{|}\bar{\theta}^{(n)}(\mathbf{r})\bigr{]}.$ (34)
Practically, the parameter values are estimated at a discrete grid of points $\mathbf{r}_{ij}$. In that case, the derivatives in Eq. (33) can be replaced by finite differences, and the values $\theta(\mathbf{r}_{ij})$ can be found by unconstrained quadratic minimization (Appendix C).
@@ II.5 Sensing method summary
2. Sensing calibration. The optimal weights $Q_{0}(x)$ and $Q_{1}(x)$ of the two detection channels are defined in Appendix A. If the experimental capabilities do not allow to perform such optimal measurement, the actual weights $Q_{0}(x)$ and $Q_{1}(x)$ are defined by the available detection setup (the profiles of the channels’ sensitivity). Based on the functions $\varepsilon(x,\theta)$, $Q_{0}(x)$, and $Q_{1}(x)$, the dependence $\zeta(\theta)$ of the compressed signal on the sensed parameter should be constructed according to Eq. (16). Then, the estimator for the $n$-th order of SOFI is described by Eq. (26).
@@
4. Data processing. For each channel $k$ and each detection position $\mathbf{r}_{ij}$ separately, the cumulant value $C_{k}^{(n)}(\mathbf{r}_{ij})=C_{t}^{(n)}[F_{k}(\mathbf{r}_{ij},t)]$ is calculated. For $n=2$, 3, 4 the expressions are provided by Eqs. (6)–(8) with the expectation values replaced by finite-sample averages. Alternatively, one can use cross-cumulants for improvement of signal-to-noise ratio as described in Appendix D. The local estimates of the parameter $\bar{\theta}^{(n)}(\mathbf{r})$ are calculated according to Eq. (26).
@@
5. Data interpolation. To account for the difference of the estimate reliability at different points, the corrected parameter distribution is constructed according to Eq. (34) with the method described in Appendix C.
III Numerical modeling
III.1 Modeled setup
To illustrate the proposed combined sensing approach, we start from its application to simulated data. The modeled object includes equidistant fluorescent probes in $8\times 8$ grid with the step $d=0.266d_{\text{R}}$ (Fig. 2) or $0.532d_{R}$ (Fig. 3), where $d_{\text{R}}$ is the Rayleigh resolution limit of the microscope for classical incoherent-light imaging. Each emitter is assigned the value $\theta_{k}=\pm 1$ of the sensed parameter $\theta$ (Figs. 2(a) and 3(a)). The modeled distribution of the parameter $\theta$ has been chosen so that sensing of features with different spatial scale can be analyzed.
The blinking of the modeled emitters is described by the telegraph process with the waiting time for switching between the bright (“on”) and dark (“off”) states distributed according to the probability densities
$p_{\text{on,off}}(\Delta t)=t_{\text{on,off}}\exp\left(-\Delta t/t_{\text{on,off}}\right),$ (35)
where $t_{\text{on}}$ and $t_{\text{off}}$ are the mean durations of continuous staying the bright and dark state respectively. The natural time unit for the modeling is the SOFI frame duration. In that units, the “on” and “off” times used for the modeling are $t_{\text{on}}=2$ frames and $t_{\text{off}}=3$ frames. The total observation time is assumed to include $1.5\times 10^{4}$ frames.
We model measurement of the fluorescence signal by a $40\times 40$ pixels array detector. The pixel size is related to the distance between the emitters as $d_{\text{pix}}:d=1:5$. The number of photons, detected from each emitter per frame (in total over the whole detection plane), obeys Poisson statistics with the expectation value $10^{3}$ photons per emitter per frame. The signal is split into two spectral channels (Fig. 1(e)) in the ratio $Y_{1}:Y_{2}=0.8:1.2$ for $\theta=+1$ and $Y_{1}:Y_{2}=0.7:1.3$ for $\theta=-1$. The signal $Y_{0}$, used for normalization in denominator of Eq. (15), is defined as the sum of both channels: $Y_{0}=Y_{1}+Y_{2}$. The SOFI images are constructed according to the approach, described in Appendix D with $\sigma=3$.
We assume that the dependence of the emitters’ response to the variations of the sensed parameter $\theta$ is linear and follows Eq. (18). According to the ratios of the channel signals, defined above, the coefficients in Eq. (19) are related as $A_{01}=0$, $A_{10}=0.375A_{00}$, and $A_{11}=0.025A_{00}$. Therefore, the estimator (19) takes the form
$\bar{\theta}=\zeta^{-1}(Z)=40Z-15.$ (36)
The source code for the discussed numerical modeling is available in Supplemental Material at [URL will be inserted by publisher].
III.2 Results
Figs. 2 and 3 show the modeling results for the cases of sub-Rayleigh and relatively well-resolved imaging with sensing respectively. One can see that, as expected, the spatial resolution of the parameter inference increases for higher orders of SOFI cumulant images. To illustrate the effect, profiles corresponding to a vertical cross-section of the maps are shown in Fig. 4. For sub-Rayleigh imaging regime (Fig. 4(a)), the fine features of the parameter distribution are not visible in the results for the first and the second order cumulants, while separate minima and maxima are already present in the profile corresponding to the fourth order cumulant (i.e., the features are resolved according to Rayleigh’s criterion). For the case shown in Fig. 4(b), the features of the parameter distribution are successfully resolved for all the considered cumulant orders. However, the accuracy of the parameter inference is much higher for the fourth order of SOFI: the mean squared error (MSE), defined as $\operatorname{MSE}(\bar{\theta})=\langle(\bar{\theta}(\mathbf{r}_{k})-\theta_{k})^{2}\rangle$ takes the values 0.40, 0.11, and 0.003 for the first, the second, and the fourth order of SOFI cumulants respectively.
Figure 2: Modeling results for sub-Rayleigh imaging with sensing. (a) The object is modeled as a grid of emitters with the step $d=0.266d_{\text{R}}$. The emitters are assigned the parameter values $\theta=+1$ (red dots) and $\theta=-1$ (blue dots). The vertical dashed line shows the cross-section used for extracting the profiles in Fig. 4. (b-d) The reconstructed distribution $\bar{\theta}^{(n)}(\mathbf{r})$ of the sensed parameter for $n=1$ (b), 2 (c), 4 (d). The color map is the same for all plots (dark blue and light yellow correspond to $\bar{\theta}=-1$ and $\bar{\theta}=+1$ respectively). The spatial coordinates $x$ and $y$ are normalized by the Rayleigh limit $d_{\text{R}}$.
Figure 3: Modeling results for relatively well-resolved imaging with sensing. The step size of emitters’ grid is $d=0.532d_{\text{R}}$. Other parameters and notations are the same as in Fig. 2.
Figure 4: Inferred parameter distributions for the vertical cross-section of the maps, shown in Figs. 2 and 3. Dot-dashed, dashed, and solid lines correspond to $n=1$, 2, and 4 respectively. Blue and red dots indicate the true values of the parameter $\theta_{k}$ ($-1$ and $+1$ respectively) for the emitters along the considered sample cross-section.
IV Model experiment
IV.1 Setup description
@@
To get closer to real-world verification of the proposed technique, we also performed the following model experiment. We used a paper screen as a model of fluorescent sample. The screen was irradiated with 13 KY-008 lasers controlled by the microcontroller board Arduino MEGA [72] and capable of independent switching between “dark” and “bright” states. The image of the screen was recorded with Thorlabs DCC1545M-GL camera with 50 mm lens (Fig. 5). Switching of lasers was synchronized to the frame rate of the camera (24 fps, approximately 42 ms per frame). To emulate the diffraction limit, the lens was placed 2 mm closer to the camera than it was necessary to get a clear picture. The resulting image of unfocused laser spots is shown in Fig. 6. Each laser spot was used for modeling a fluorescent particle (temperature nanosensor).
@@
Figure 5: The scheme of the model experiment. The lasers controlled with Arduino microcontroller board irradiate the paper screen. The light scattered by the screen is collected with lens to the Thorlabs camera.
To model fluorescent particles with different temperature, the lasers were divided into “hot” ($\theta=+1$) and “cold” ($\theta=-1$) ones. The parameter value $\theta$ (normalized “temperature”) was encoded in the “relaxation time” — the time between “excitation pulse” and “fluorescent emission”. To denote “excitation pulse” during data recording, all lasers blinked twice (i.e., a series of 2 one-frame flashes of lasers separated by 1 frame of darkness was produced). The “fluorescent emission” was 1-frame flashing of a laser. For the “hot” lasers, the “relaxation time” was 5 frames with standard deviation of 2 frames, while the “cold” lasers had “relaxation time” of 10 frames with the same standard deviation.
@@
Figure 6: Image of blurred laser spots on the paper screen for the lens being placed 2 mm closer to the camera than it is necessary for a clear image. Numbers are used for indexing the laser beams during further analysis.
IV.2 Data acquisition
The data acquisition was based on 3 time scales (Fig. 7):
  • Camera frame (42 ms) represented the natural elementary unit of time.
  • “Excitation cycle” lasted 20 camera frames and included an “excitation pulse” and “fluorescent emission” of each active laser.
  • SOFI frame included 20 “excitation cycles” (400 camera frames) and determined blinking rate of the modeled fluorescent particles.
@@ IV.2 Data acquisition
For each SOFI frame, each laser was randomly assigned a “bright” (active) or “dark” (inactive) state to model stochastic blinking. The probabilities of the two states were equal to 1/2, and the sampling was performed independently for all lasers and all SOFI frames. For each “excitation cycle” withing a SOFI frame, each laser being in a “bright” state was randomly assigned its relaxation time (delay of “fluorescent emission”). The delay was sampled according to normal distribution with the mean determined by the model temperature of the laser (“hot” or “cold”). The random sampling was performed by the Arduino microcontroller. Then, the microcontroller produced an “excitation pulse” and made each active laser to flash after the time delay assigned to it. After modeling “fluorescent emission” of all the active lasers withing the current cycle (when all the lasers are switched off), the microcontroller waited for 1 camera frame, and repeated the cycle. In the end, we got a video containing intensity distribution in time and space. The total acquisition process included $2\times 10^{5}$ camera frames ($10^{4}$ excitation cycles, and 500 SOFI frames).
IV.3 Data processing and results
@@
First, the recorded video was divided into frames with ffmpeg program [73] and processed with Python script. The script calculated total intensity for each camera frame and identified “excitation pulses” according to specific alternation of high and low intensity of frames. The frames designating the “excitation pulse” were excluded from further calculations. The resulting series of “excitation cycles” was split into groups of 20 cycles, each group forming a SOFI frame. According to the notations, introduced in Section II, the SOFI frames are indexed by $t=1$, …, $N_{\text{f}}$.
@@
The signal for each “excitation cycle” was split into two channels and accumulated pixelwise. Each SOFI frame accumulates the signals from the “excitation cycles” included in that frame. The “hot” signal $Y_{1}(i,j;t)$ is the pixelwise sum of intensities for the 1st - 7th camera frames after the $t$-th “excitation pulse” accumulated for 20 cycles, where indices $i$ and $j$ enumerate pixel positions. The “cold” signal $Y_{2}(i,j;t)$ is defined in the same way starting from the 8th camera frame after the “excitation pulse” (Fig. 7).
@@
Figure 7: Data acquisition timeline. (a) A SOFI frame consist of 20 “excitation cycles”. (b) An example of “excitation cycle” timeline, where two pulses appear in the “hot” time window and two pulses - in the “cold” time window.
@@
The intensity (mean signal value) and the 2nd, 3rd, and 4th order cumulants for the two channels where defined according to Eqs. (6)–(9). Here, only zero time lag autocumulants were used since the signal did not suffer from significant shot noise. The obtained intensity and cumulant images are shown in Fig. 8. For visualization purposes, the “cold” and “hot” channels were assigned to the blue and red channels of RGB palette respectively. Since the cumulants of the 3rd and the 4th order can be negative, the absolute values of the cumulants were used to create the image. One can see that the intensity distribution (Fig. 8(a)) is blurred like original unfocused image (Fig. 6(b)). However, it is possible to distinguish the light of “hot” and “cold” lasers because of different relaxation times. The light spots are more visible in the 2nd order cumulant image (Fig. 8(b)), but the effect of blurring is still strong. The 3rd and the 4th order cumulant images (Fig. 8(c,d)) have the best quality and allow one to distinguish clearly spots of “hot” and “cold” lasers. To summarize, the experiment is in a good agreement with the theory, and the calculation of cumulants of intensity fluctuations allows one to visualize the particles with different relaxation times.
@@
IV.4 Reconstruction of local “temperature”
@@
To fit the above-described model experiment into the theoretical framework, discussed in Section II, we assign the normalized dimensionless “temperature” equal to $\theta=-1$ to “cold” lasers and $\theta=+1$ to “hot” lasers. The emission profile variable $x$ corresponds to the camera frame number within the “excitation cycle” in the considered case. For such definitions, the model emission profile is described as
$\varepsilon(x,\theta)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{(x-7.5+2.5\theta)^{2}}{2\sigma^{2}}\right),$ (37)
where $\sigma=2$.
A simple threshold filter is used to split the two detection channels (see Fig. 1(f)):
$Y_{1}=\sum_{x\leq 7}\varepsilon(x,\theta),\quad Y_{2}=\sum_{x\geq 8}\varepsilon(x,\theta),\quad Y_{0}=Y_{1}+Y_{2}.$ (38)
The function $\zeta(\theta)$, calculated according to Eq. (16), is plotted in Fig. 9. The estimators $\bar{\theta}^{(n)}(\mathbf{r})$ are defined by Eq. (26). Further, the reconstructed distribution can be refined using the method outlined in the Section II.4 and Appendix C.
Figure 9: Relation between the compressed signal $Z=\zeta(\theta)$ and the sensed parameter $\theta$ (normalized dimensionless “temperature”) for the model experiment, calculated according to Eqs. (16), (37), and (38).
@@
The results are presented in Figs. 10 and 11. As predicted, usage of higher-order cumulants enhances spatial resolution and enables analysis of finer features of the sample. According to the standard performance of SOFI, using fourth order cumulant instead of intensity leads to approximately twice narrower efficient PSF.
Figure 10: Reconstructed distribution of the model normalized temperature $\theta$ for analysis of intensity (a), the 2nd (b), 3rd (c), and 4th (d) order cumulants. Black lines indicate the border of the region where non-zero signal was collected; outside that region the temperature values are effectively extrapolated by the approach presented in Subsection II.4. Black dots shows the central positions of the laser spots (positions of the modeled fluorescent emitters).
Figure 11: Cross-section of the reconstructed distributions of the model normalized temperature $\theta$, shown in Fig. 10, for the vertical line passing through the emitter no. 3. Black markers shows true values of the parameter $\theta$ for the emitters no. 1, 3, 5, and 9 adjacent to the considered cross-section.
IV.5 Discussion
To get quantitative estimates of the proposed approach performance, we compared the estimated parameter values $\bar{\theta}_{k}=\bar{\theta}(\mathbf{r}_{k})$ at the positions of the modeled fluorescent particles (centers of the laser spots) with the true values $\theta_{k}=\pm 1$ (Fig. 12(a)). Fig. 12(b) shows the dependence of the mean squared error (MSE), defined as $\operatorname{MSE}(\theta)=\langle(\bar{\theta}_{k}-\theta_{k})^{2}\rangle_{k}$, on the cumulant order $n$. The results indicate that the least accurate estimate corresponds to the blurred distribution obtained for the traditional intensity-based approach. The estimation accuracy is saturated for $n=3$, while for $n=4$ the error even increases slightly because of stronger artifacts and noise enhancement typical for higher-order cumulants. Such saturation is explained by the specific profile of blurring (containing a plateau, as one can see from Fig. 11) — see Appendix E for details.
Figure 12: Performance of the proposed approach. (a) Comparison of estimated parameter values $\bar{\theta}_{k}=\bar{\theta}(\mathbf{r})_{k})$ at the positions of the modeled fluorescent particles (centers of the laser spots) with the true values $\theta_{k}=\pm 1$ for different orders of cumulants used for the analysis (blue circles, yellow squares, green diamond makers, and red stars correspond to $n=1$, 2, 3, and 4 respectively). (b) Mean squared error of the estimation as the function of the cumulant order $n$.
V Conclusions
We proposed a simple and efficient technique capable of combining sensing with super-resolution imaging, based on non-trivial temporal statistics typical for blinking fluorescent nanoemitters. The approach is universal since the dependence of the fluorophores response on the characteristic to be reconstructed can take different form: changes of the spectrum, fluorescence lifetime, etc. It is shown that splitting the fluorescence signal into just two detection channels at each spatial detection point is sufficient for reliable reconstruction of the local characteristic distribution.
The conducted numerical simulations and model experiment proof soundness and efficiency of the approach and can be used as a guidance for the real-world experiments with, for example, living cells and structured fluorescent nanodiamonds [74] or blinking graphene quantum dots ensembles [75]. In agreement with the theoretical predictions, usage of cumulants instead of emission intensity enhances sharpness of the reconstructed parameter distribution, enables analysis of smaller features of the sample, and ensures higher accuracy of the estimation.
Proposed methodology will be useful for local measurements of living cells’ physical parameters, such as temperature, pH, electric field gradient, presence of reactive oxygen spices, i.e. the characteristics which could be recognized as spectral changes of appropriate fluorophore delivered targetly to the wanted cell or even organelle. In its turn, this advanced local sensing tool could give an important insights into fundamental life science, which still lacks instruments capable of monitoring the cellular balance with nanometer spatial resolution to guarantee measurements consistency for that dynamic system.